# MAGNET: an open-source library for mesh agglomeration by Graph Neural Networks

### v1

@@ 1 Introduction
In this work, we introduce MAGNET (Mesh Agglomeration by Graph Neural Network), an open-source Python library, released under the GNU Lesser General Public License, version 3 (LGPLv3). MAGNET provides a simple framework for mesh agglomeration in both two- and three-dimensions, it allows experimenting with different neural network architectures, training them, and comparing their performance to state-of-the-art methods like METIS [23] and k-means on standard quality metrics. MAGNET can also be easily integrated with other software, in particular, it already interfaces with lymph [27], an open-source Matlab library for PolyDG methods: this allows an assessment of the agglomerated meshes by solving differential problems on it. Our approach extends the results presented in [28, 2] and consists of re-framing the problem of mesh agglomeration as a graph partitioning problem by exploiting the dual mesh. A key innovation of the present work is the adoption of the reinforcement learning (RL) approach introduced in [29] to train the model, offering a novel paradigm to adaptively optimize graph-based partitioning in heterogeneous domains. One of the main advantages of GNN-based algorithms is the possibility to incorporate the physical properties of the computational domain during mesh agglomeration and thus taking into account heterogeneities of the computational physical domain, embedded microstructures, and/or interior barriers, and preserve them during the agglomeration process. We remark that our library is independent of both the PDE and the polytopal method under investigation, as it solely relies on the underlying grid agglomeration procedure, making it adaptable in principle for integration into any polytopal code.
@@ 2.1 Mesh Agglomeration and a Graph Partitioning Problem
where $\text{deg}(v)=|(v,v_{i})\in\mathcal{E},\forall\ v_{i}\in\mathcal{V}|$ is the degree of node $v$, and ||$\cdot$|| ⋅ | indicates the cardinality of the set. The definition of cut can be extended to a generic partition $(\mathcal{S}_{1},\dots,\mathcal{S}_{M})$, where $M\in\mathbb{N}$ is the total number of subsets in the partition:
@@ 2.2.1 The SAGE-Base Model
The standard GNN model implemented in MAGNET, that we will call SAGE-Base in the following, employs four SAGE convolutional layers (sample and aggregate [36]) followed by four linear, or dense, layers; the hyperbolic tangent is used after each layer to keep the output in the range $[-1,1]$ and facilitate learning. A final softmax layer ensures that the GNN output is a probability distribution over the two classes for each node. We use as node features the coordinates of the centroids of each mesh element and the measure of each element, so that we have a total of four node features (three in the two-dimensional case). Before feeding them to the GNN, an additional normalization step, consisting of normalizing coordinates to zero mean and unit variance and rotating the mesh so that the widest direction is aligned with the x-axis, is performed to reduce variability in input. Since graphs are dimensionless objects, the approach is the same for polygonal and polyhedral meshes: only the number of input features and the normalization procedure have to be adapted between the two. The two-dimensional version of this model has been trained on a set of 800 meshes comprising structured quadrilaterals, structured triangles, random Delaunay triangulations, and random Voronoi tessellations in equal measure. The training has been performed using the Adam optimizer [37] for 300 epochs with learning rate $\gamma=10^{-5}$, weight decay parameter $\lambda=10^{-5}$ and batch size 4. The three-dimensional version has a greater number of parameters to account for the increased variability of the three-dimensional case, doubling the size of SAGE and linear layers from 64 and 32 to 128 and 64 units, respectively. The training dataset comprised 400 tetrahedral meshes, of which 100 were of the unit cube and 300 were of randomly generated portions of it (see Figure 4g); other changes include using a learning rate of $\lambda=10^{-4}$ and lengthening training to 400 epochs. In both cases, meshes have been randomly rotated during training to artificially augment the dataset.
@@ 2.2.2 The SAGE-Heterogeneous Model
$\frac{a}{|\mathcal{V}|}\sum_{i=1}^{N}\sum_{j=1}^{2}\big{(}\mathbf{P}\odot\mathbf{Y}\big{)}_{ij},$
where $\mathbf{P}\in\mathbb{R}^{N\times 2}$ is a matrix that contains the physical parameter $p$ and its complementary $1-p$ on each row, $\odot$ denotes element-wise multiplication, $|\mathcal{V}|$ is the number of nodes in the graph and $a$ is a suitable coefficient chosen by performing a hyperparameter search; the term is normalized with respect to graph size to make it scale invariant. While this model works only when there are at most two different physical parameters, this approach could be extended to an arbitrary number of different physical groups by adding as node features a one-hot encoding of the various heterogeneous parts; however, this would require a different GNN for each total number of physical groups. The obtained model is referred to as SAGE-Heterogeneous in the code. The bisection algorithm is the same since the output matrix $\mathbf{Y}$ has the same meaning. The SAGE-Heterogeneous model included in MAGNET has the same architecture as the SAGE-Base model and was trained using mostly the same hyperparameters. The two-dimensional version has been trained for 200 epochs on a dataset of 600 meshes of the unit square, of which 200 are homogeneous, 200 are divided into two parts along a line, and 200 have 6-12 circular inclusions. The three-dimensional version has been trained on a set of 400 meshes of the unit cube divided into one to four physically heterogeneous parts each for 150 epochs.
@@ 2.3.1 The Reinforcement Learning Partitioner
[…]
@@ 2.4 Other Agglomeration Strategies Available in MAGNET
In addition to the GNN strategies described in the previous Section 2.2 and Section 2.3, MAGNET also includes both METIS and k-means as possible agglomeration approaches. METIS [23] is a widely used library for computing k-way partitions of large irregular graphs based on a multilevel procedure. It uses simple and efficient algorithms (like greedy approaches) for partitioning the coarsest graph, while the refinement is performed based on local heuristics like KL and FM, while respecting balancing constraints. To agglomerate a mesh, METIS is run with the requirement of creating connected subgraphs; also, we use as node weights the volumes of the corresponding cells so that the balance constraint in the partitioning algorithm is imposed in terms of geometric size, which is more relevant to our application than simple node cardinality. K-means [42] is a quantization algorithm that clusters a set of points in $\mathbb{R}^{N}$ into $k$ sets by minimizing the squared Euclidean distances of each point to the centroid of its group, i.e. by minimizing the within-group variance. While it cannot be applied directly to graphs because they are non-Euclidean data, k-means can be used with node embedding techniques to map each node onto a feature space, producing an effective clustering of the nodes. Indeed, instead of applying k-means to the mesh graph, we use it to directly cluster the cells by using as features the centroid coordinates of the mesh elements, producing fairly rounded agglomerated elements of similar size. In this way, we are neglecting the mesh connectivity, but this is usually not an issue because for meshes with similarly sized elements, centroid coordinates are strongly correlated with two cells being adjacent.
@@ 3.1 Code Structure
MAGNET is organized in the following sub-modules:
@@
The package additionally includes the agglomerate.py script, which can be called from the command line to agglomerate a single mesh, and is also used for communication with the lymph library (see Section 5). Finally, the models folder contains some pre-trained GNN models according to the specifications detailed in the previous sections.
@@ 3.1.2 The IO Sub-module
7:$\mathbf{A}_{ij}\leftarrow 0,i,j=1,\dots,N$triangleright Initialize adjacency matrix as sparse empty matrix
@@ 3.2 Agglomeration Algorithms
  • multilevel recursive bisection: the mesh graph is initially coarsened using a greedy heavy edge matching algorithm [25], then bisected; the coarse partition is successively projected unto the finer graph and refined (using, for example, the RL Refiner of Section 2.3.2); this process is recursively repeated to obtain the desired partition. See _absaggmodels.py line 382 for details.
@@
[…]
@@
2:$\texttt{parts}\leftarrow[\mathcal{G}]$
3:$\texttt{output}\leftarrow[]$
@@
7:         if diam$(\mathcal{S})$ > $\hat{h}$ then
8:              $\mathcal{S}_{1},\mathcal{S}_{2}\leftarrow$ bisect($\mathcal{S}$) $\triangleright$ If the element is too "large", bisect it
@@
11:              output.append$(\mathcal{S})$triangleright Otherwise, insert it in the final output
@@ 3.5 Quality Metrics
    This quantity can take any positive value, so to have a metric between zero and one, we consider instead:
@@ 4.3 Test Case 2: Two-Dimensional Domain with Inclusions
We consider a set of four triangular meshes of the unit square with an increasing number of circular inclusions that represent microstructures in the underlying domain, with respectively 4, 16, 36, and 64 inclusions each. The radii of the circles have been chosen so that they always cover 15% of the total area. We agglomerate them both with the SAGE-Heterogeneous model and by using segregated mode with both METIS and SAGE-Base using a target size equal to 3/4 of the circle diameter, so that the agglomerated elements have a size comparable to that of the inclusions. The results are reported in Figure 8.
@@
In this case, the SAGE-Heterogeneous model correctly separates the inclusions; however, since from the second bisection onward only one of the two physical groups appears, k-means is called instead.
@@ 4.6 Test Case 5: Three-Dimensional Brain Geometry
To demonstrate the generalization capabilities of the models in the three-dimensional case, we consider once again a mesh coming from an MRI scan of a human brain, shown in Figure 11. We agglomerate the mesh using four different models (METIS, k-means, SAGE-Base, and SAGE-Base as coarse partitioner in a multilevel framework with RL Refiner). For the GNN models, we used number of refinements mode with nref=8, while for the other two direct k-way with k=256 was used. In Figure 12 we report the agglomerated meshes and their exploded view, together with their quality metrics box plots. Most of the same considerations of the two-dimensional case hold, with some differences: due to the highly corrugated surface of the brain, k-means and SAGE-Base struggle to create fully connected elements, often leaving a few tetrahedra separated from the rest. As a consequence, the corresponding elements will be much smaller, leading to a lower uniformity factor. This issue can be alleviated in the k-means case by properly choosing a higher $k$, while not much can be done for SAGE-Base since the problem arises from the very first bisection. On the other hand, the multilevel approach with the RL refinement mostly solves the issue, improving the quality of the partition.
@@ 5 Integration with the lymph Library
5.1 Verification Test 1: Poisson Problem on a Brain Slice
@@
We discretize problem (8) with a PolyDG method as shown in [27]. For simplicity, we consider as analytical solution $u(\textbf{x})=\sin(2\pi\textbf{x}_{1})\cos(2\pi\textbf{x}_{2})$ and compute the right-hand side $f$ and boundary condition $g$ accordingly. We solve this problem on the brain slice we used in Section 4.2 rescaled to fit in $(0,1)^{2}$. The mesh is agglomerated from Matlab by calling the Agglomerate.m function as reported in Listing LABEL:lst:poissonbrain.

### v2

@@ 1 Introduction
In this work, we introduce MAGNET (Mesh Agglomeration by Graph Neural Network), an open-source Python library, released under the GNU Lesser General Public License, version 3 (LGPLv3). MAGNET provides a simple framework for mesh agglomeration in both two- and three-dimensions, it allows to experiment with different neural network architectures, training them, and comparing their performance to state-of-the-art methods like METIS [23] and k-means on standard quality metrics. MAGNET can also be easily integrated with other software, in particular, it already interfaces with lymph [27], an open-source Matlab library for PolyDG methods: this allows an assessment of the agglomerated meshes by solving differential problems on it. Our approach extends the results presented in [28, 2] and consists of re-framing the problem of mesh agglomeration as a graph partitioning problem by exploiting the dual mesh. A key innovation of the present work is the adoption of the reinforcement learning (RL) approach introduced in [29] to train the model, offering a novel paradigm to adaptively optimize graph-based partitioning in heterogeneous domains. One of the main advantages of GNN-based algorithms is the possibility to incorporate the physical properties of the computational domain during mesh agglomeration and thus taking into account heterogeneities of the computational physical domain, embedded microstructures, and/or interior barriers, and preserve them during the agglomeration process. We remark that our library is independent of both the PDE and the polytopal method under investigation, as it solely relies on the underlying grid agglomeration procedure, making it adaptable in principle for integration into any polytopal code.
@@ 2.1 Mesh Agglomeration and a Graph Partitioning Problem
where $\text{deg}(v)=|(v,v_{i})\in\mathcal{E},\forall\ v_{i}\in\mathcal{V}|$ is the degree of node $v$, and $|\cdot|$ indicates the cardinality of the set. The definition of cut can be extended to a generic partition $(\mathcal{S}_{1},\dots,\mathcal{S}_{M})$, where $M\in\mathbb{N}$ is the total number of subsets in the partition:
@@ 2.2.1 The SAGE-Base Model
The standard GNN model implemented in MAGNET, that we will call SAGE-Base in the following, employs four SAGE convolutional layers (sample and aggregate [36]) followed by four linear, or dense, layers; the hyperbolic tangent is used after each layer to keep the output in the range [1,1][-1,1] and facilitate learning. A final softmax layer ensures that the GNN output is a probability distribution over the two classes for each node. We use as node features the coordinates of the centroids of each mesh element and the measure of each element, so that we have a total of four node features (three in the two-dimensional case). Before feeding them to the GNN, an additional normalization step, consisting of normalizing coordinates to zero mean and unit variance and rotating the mesh so that the widest direction is aligned with the x-axis, is performed to reduce variability in input. Since graphs are dimensionless objects, the approach is the same for polygonal and polyhedral meshes: only the number of input features and the normalization procedure have to be adapted between the two. The two-dimensional version of this model has been trained on a set of 800 meshes comprising structured quadrilaterals, structured triangles, random Delaunay triangulations, and random Voronoi tessellations in equal measure. The training has been performed using the Adam optimizer [37] for 300 epochs with learning rate $\gamma=10^{-5}$, weight decay parameter $\lambda=10^{-5}$ and batch size 4. The three-dimensional version has a greater number of parameters to account for the increased variability of the three-dimensional case, doubling the size of SAGE and linear layers from 64 and 32 to 128 and 64 units, respectively. The training dataset comprised 400 tetrahedral meshes, of which 100 were of the unit cube and 300 were of randomly generated portions of it (see Figure 4g); other changes include using a learning rate of $\lambda=10^{-4}$ and lengthening training to 400 epochs. In both cases, meshes have been randomly rotated during training to artificially augment the dataset.
@@ 2.2.2 The SAGE-Heterogeneous Model
$\frac{a}{|\mathcal{V}|}\sum_{i=1}^{N}\sum_{j=1}^{2}\big(\mathbf{P}\odot\mathbf{Y}\big)_{ij},$ (2)
where $\mathbf{P}\in\mathbb{R}^{N\times 2}$ is a matrix that contains the physical parameter $p$ and its complementary 1p1-p on each row, $\odot$ denotes element-wise multiplication, $|\mathcal{V}|$ is the number of nodes in the graph and $a$ is a suitable coefficient chosen by performing a hyperparameter search; the term is normalized with respect to graph size to make it scale invariant. While this model works only when there are at most two different physical parameters, this approach could be extended to an arbitrary number of different physical groups by adding as node features a one-hot encoding of the various heterogeneous parts; however, this would require a different GNN for each total number of physical groups. The obtained model is referred to as SAGE-Heterogeneous in the code. The bisection algorithm is the same since the output matrix $\mathbf{Y}$ has the same meaning. The SAGE-Heterogeneous model included in MAGNET has the same architecture as the SAGE-Base model and was trained using mostly the same hyperparameters. The two-dimensional version has been trained for 200 epochs on a dataset of 600 meshes of the unit square, of which 200 are homogeneous, 200 are divided into two parts along a line, and 200 have 6-12 circular inclusions. The three-dimensional version has been trained on a set of 400 meshes of the unit cube divided into one to four physically heterogeneous parts each for 150 epochs.
@@ 2.3.1 The Reinforcement Learning Partitioner
Refer $\rightarrowRefer to caption\rightarrow\dots\rightarrowRefer$ to caption
@@ 2.4 Other Agglomeration Strategies Available in MAGNET
In addition to the GNN strategies described in the previous Section 2.2 and Section 2.3, MAGNET also includes both METIS and k-means as possible agglomeration approaches. METIS [23] is a widely used library for computing k-way partitions of large irregular graphs based on a multilevel procedure. It uses simple and efficient algorithms (like greedy approaches) for partitioning the coarsest graph, while the refinement is performed based on local heuristics like KL and FM, while respecting balancing constraints. To agglomerate a mesh, METIS is run with the requirement of creating connected subgraphs; also, we use as node weights the volumes of the corresponding cells so that the balance constraint in the partitioning algorithm is imposed in terms of geometric size, which is more relevant to our application than simple node cardinality.
K-means [42] is a quantization algorithm that clusters a set of points in $\mathbb{R}^{N}$ into $k$ sets by minimizing the squared Euclidean distances of each point to the centroid of its group, i.e. by minimizing the within-group variance. While it cannot be applied directly to graphs because they are non-Euclidean data, k-means can be used with node embedding techniques to map each node onto a feature space, producing an effective clustering of the nodes. Indeed, instead of applying k-means to the mesh graph, we use it to directly cluster the cells by using as features the centroid coordinates of the mesh elements, producing fairly rounded agglomerated elements of similar size. In this way, we are neglecting the mesh connectivity, but this is usually not an issue because for meshes with similarly sized elements, centroid coordinates are strongly correlated with two cells being adjacent.
@@ 3.1 Code Structure
MAGNET is organized as reported in Figure 3, namely it is composed by the following sub-modules:
@@
The package additionally includes the lymphcomm.py script, which can be called from the command line to agglomerate a single mesh, and is also used for communication with the lymph library (see Section 5). Finally, the models folder contains some pre-trained GNN models according to the specifications detailed in the previous sections.
@@ 3.1.2 The IO Sub-module
[…]
@@ 3.2 Agglomeration Algorithms
  • multilevel recursive bisection: the mesh graph is initially coarsened using a greedy heavy edge matching algorithm [25], then bisected; the coarse partition is successively projected unto the finer graph and refined (using, for example, the RL Refiner of Section 2.3.2); this process is recursively repeated to obtain the desired partition. See _absaggmodels.py line 395 for details.
@@
Input: $\mathcal{G},\texttt{mult\_factor}$: graph to be partitioned, multiplicative factor
@@
4:while $\texttt{parts}\ \textbf{is not}\ \emptyset$ do
5:  $\texttt{new\_parts}\leftarrow[]$
@@
8:     $\mathcal{S}_{1},\mathcal{S}_{2}\leftarrow$ bisect($\mathcal{S}$) $\triangleright$ If the element is too “large", bisect it
9:     $\texttt{new\_parts}\leftarrow\texttt{new\_parts}+[\mathcal{S}_{1},\mathcal{S}_{2}]$
@@
[…]
@@ 3.5 Quality Metrics
    This quantity can take any positive value, so to have a metric between zero and one we consider instead:
@@
  • Heterogeneity Preservation (HP): the relative presence of different physical tags $p_{i}=0,1$ in an agglomerated cell $P$ of $n_{P}$ fine cells.
    $\text{HP}(P)=\max\{\bar{p},1-\bar{p}\},\quad\bar{p}=\frac{1}{n_{P}}\sum_{i}^{n_{P}}p_{i}$ (9)
    A value of one denotes that all elements have the same physical tag, while lower values correspond to a less homogeneous distribution across the cells.
@@ 4.1 GNN Training
4.2 Computational Cost
In this section, we explore the advantages of the GNN-based agglomeration from a point of view of computational cost. In particular, we provide quantitative measurements of the wall-clock time for METIS, k-means, and GNN-based agglomeration. For the GNN, we compare the wall-clock time for both a CPU-based and a GPU-based run. In order to provide a comparison that is as fair as possible, we use the resources offered for free in Google Colab, namely an Intel(R) Xeon(R) CPU @ 2.20GHz and an Nvidia T4 GPU. In Figure 6, we provide an in-depth analysis of the computational cost depending on the space dimension (2D/3D) and the number of refinements. If we compare the k-way and the bisect variants, we see that the k-way algorithm is much cheaper than the bisect version of METIS. This holds in both 2D and 3D since METIS considers only the dual graph and is thus space independent. On the other hand, k-means scales linearly with respect to the space dimension, the number of cells, and the number of clusters. Hence, the k-way version of k-means performs worse than the bisect version as we refine the mesh (namely, the number of cells and the number of refinements increase). This is particularly evident in the case nref=8 where we can see that the k-way k-means has a much steeper slope compared to the other methods. When running the GNN method on a CPU, the cost is asymptotically comparable to the bisect version of METIS. On the other hand, the GPU version, while having a larger fixed cost until around $10^{4}$ cells, is the cheaper method for larger meshes. These results strongly suggest that the GPU hardware is essential to fully unlock the scalability benefits of GNN-based agglomeration, making it a cheaper alternative to traditional methods for medium to large-scale problems.
Figure 6: Comparison of the computational cost of the algorithms present in MAGNET for the agglomeration of 2D meshes (top) and 3D meshes (bottom).
@@ 4.4 Test Case 2: Two-Dimensional Domain with Inclusions
We consider a set of four triangular meshes of the unit square with an increasing number of circular inclusions that represent microstructures in the underlying domain, with respectively 4, 16, 36, and 64 inclusions each. The radii of the circles have been chosen so that they always cover 15% of the total area. We agglomerate them both with the SAGE-Heterogeneous model and by using segregated mode with both METIS and SAGE-Base using a target size equal to 3/4 of the circle diameter, so that the agglomerated elements have a size comparable to that of the inclusions. The results are reported in Figure 9. In this case, the SAGE-Heterogeneous model correctly separates the inclusions; however, since from the second bisection onward only one of the two physical groups appears, k-means is called instead.
A sensitivity analysis on the parameter $a$ of Eq. (2) is provided in Figure 10. It shows that most metrics (Sphericity, Uniformity) remain largely unaffected, while Circle Ratio and Volume Difference display higher variability. Heterogeneity Preservation is highly sensitive, with values below 0.95 indicating poor performance. In particular, a value of $a>0.5$ seems to be the minimum to guarantee consistent results in heterogeneous agglomeration.
@@
Figure 10: Test Case 2: sensitivity analysis for the parameter $a$ of the loss function Eq. (2). On each row, the model is tested on a different mesh containing 4, 16, 36, 64 circular inclusions. The four test meshes are the ones considered also in Figure 9.
@@ 4.7 Test Case 5: Three-Dimensional Brain Geometry
To demonstrate the generalization capabilities of the models in the three-dimensional case, we consider once again a mesh coming from an MRI scan of a human brain, shown in Figure 13. We agglomerate the mesh using four different models (METIS, k-means, SAGE-Base, and SAGE-Base as coarse partitioner in a multilevel framework with RL Refiner). For the GNN models, we used number of refinements mode with nref=8, while for the other two direct k-way with k=256 was used. In Figure 14 we report the agglomerated meshes and their exploded view, together with their quality metrics box plots. Most of the same considerations of the two-dimensional case hold, with some differences: due to the highly corrugated surface of the brain, k-means and SAGE-Base struggle to create fully connected elements, often leaving a few tetrahedra separated from the rest. As a consequence, the corresponding elements will be much smaller, leading to a lower uniformity factor. This issue can be alleviated in the k-means case by properly choosing a higher $k$, while not much can be done for SAGE-Base since the problem arises from the very first bisection. The occurrence of disconnected elements is primarily inherent to the architecture of GNNs. In particular, the message-passing layers preserve knowledge of the topology only within a limited neighborhood of each node in the dual graph. Extending the receptive field to capture more global connectivity would mitigate this issue, but at the expense of significantly increasing computational cost. On the other hand, the multilevel approach with the RL refinement mostly solves the issue, improving the quality of the partition.
@@ 5 Integration with the lymph Library
5.1 Verification Test 1: Poisson Problem
@@
We discretize problem (10) with a PolyDG method as shown in [27]. For simplicity, we consider as analytical solution $u(\textbf{x})=\sin(2\pi\textbf{x}_{1})\cos(2\pi\textbf{x}_{2})$ and compute the right-hand side $f$ and boundary condition $g$ accordingly. We consider two different geometries for $\Omega$. First, we consider the unit square $\Omega=(0,1)^{2}$ as a baseline benchmark and then the brain slice we used in Section 4.3 as validation for real-world geometries.
5.1.1 Poisson Problem in the Unit Square
We consider a structured triangular mesh of the unit square with 1024 elements per side, of the kind shown in Figure 4 (a). We agglomerate the mesh with four different methods (METIS (k-way and bisect), k-means (bisect) and SAGE-Base); we skip the k-way version of k-means due to the large computational cost. In Figure 17 we show that we obtain the expected convergence rate for all the methods. However, the higher quality mesh obtained with SAGE and k-means achieve smaller errors with the same number of degrees of freedom, showing that the improved mesh quality directly translates into better accuracy for the same number of degrees of freedom.
5.1.2 Poisson Problem in a Brain Slice
We solve this problem on the brain slice we used in Section 4.3 rescaled to fit in $(0,1)^{2}$. The mesh is agglomerated from Matlab by calling the Agglomerate.m function as reported in Listing LABEL:lst:poissonbrain.